OpenStudy (anonymous):

Please Help! :) Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

Correct me if I'm wrong, but I believe that this is how I would solve the first problem. T+B = 56 B = 8 + 2T T + 8 + 2t = 56 3T + 8 = 56 3T = 48 T = 16 16 * 2 = 32 32 + 8 = 40. Tony is 16, and Belinda is 40. :)

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

Ok, thank you, Can you help me solve the other two please? =)

OpenStudy (anonymous):

2. A+S=150 A=150-S 45A+28.5S=4770 45(150-S)+28.5S=4770 6750-45S+28.5S=4770 -16.5S=-1980 S=120 A=150-120 A=30 3. D+Q=46 D=46-Q .1D+.25Q=7 .1(46-Q)+.25Q=7 4.6-.1Q+.25Q=7 .15Q=2.4 Q=16 D=46-16 D=30

OpenStudy (anonymous):

2. A=30 S=120 3. D=30 Q=16

OpenStudy (anonymous):

NIALL!